OpenStudy (anonymous):

Solve 7.74x + 3 = 10.2

hero (hero):

First, subtract 3 from both sides. Let me know what you get.

OpenStudy (anonymous):

And then after what hero said, divide by 7.74

hero (hero):

One step at a time @Veelution

OpenStudy (anonymous):

7.74x= 7.20

hero (hero):

Great. Now, as Veelution said, divide both sides by 7.74

OpenStudy (anonymous):

oh wait ; i wrote the equation wrong : / It's actually 7.7^4x+3=10.2

hero (hero):

Wow, that greatly changes the outcome

hero (hero):

Take logs of both sides, then

OpenStudy (anonymous):

how in the would do i do that ?

OpenStudy (anonymous):

i'm sorry , but i reall suck at math -_-

hero (hero):

\(\log(7.7)^{4x}= \log(7.20)\)

OpenStudy (anonymous):

i know the answer

hero (hero):

@Snapbacklive, let me remind you again that giving answers is not helpful. Help students get to the answer, not give them.

OpenStudy (anonymous):

i know but i didn't right

hero (hero):

No, but you probably wanted to.

OpenStudy (anonymous):

@Hero but i can't put the x into the equation in my calculator . so what do i do with it ? & i got the log of 7.20 .

OpenStudy (anonymous):

x is nothing

hero (hero):

You didn't have to necessarily use your calculator just yet. Usually, you want to wait until you have completely isolated x first.

OpenStudy (anonymous):

it's just a sub

OpenStudy (anonymous):

ahh , all of this is beyond confusing -_-

hero (hero):

But anyway, here's the rule you're looking for @kianalauntra \(\log(b)^a = a \log(b)\)

OpenStudy (anonymous):

well it depends

OpenStudy (anonymous):

i got that . . .

hero (hero):

Okay, here's what you should have at this point: \(4x \log(7.7) = \log(7.20)\)

hero (hero):

Do you have an idea of what the next step might be?

OpenStudy (anonymous):

do i get the logs of both ?

hero (hero):

We should wait until the end before computing logs. For the next step, you want to divide both sides by \(\log(7.7)\). Remember we're trying to isolate x first.

hero (hero):

Let me know what you get after dividing both sides by \(\log(7.7)\)

OpenStudy (anonymous):

so it would be 4x=log(7.20)/log(7.7) ?

hero (hero):

Yes, very good. Now, divide both sides by 4.

OpenStudy (anonymous):

wait , how would i divide both sides by 4 if im already dividing the two logs ?

hero (hero):

Multiply both sides by 1/4. Do it that way and let me know what you get.

OpenStudy (anonymous):

x=log(7.7)/log(7.20)*0.25

OpenStudy (anonymous):

i meant log(7.20)/log(7.7)

hero (hero):

I would have written it like this just to be safe \[\large x = \frac{\log(7.20)}{4 \log(7.7)}\]

hero (hero):

Now, you can use your calculator to figure out an approximate value of x

OpenStudy (anonymous):

omigod , thank you so much !

OpenStudy (anonymous):

so x=.86/3.55 = .24 ?

hero (hero):

Yes, correct, although anytime you have a decimal answer like that, you should probably write \(x \approx .24\) since it is a decimal approximation.

hero (hero):

Great job :D

OpenStudy (anonymous):

thank you <333 .